# Gold after Randomized Sand: Model-X Split Knockoffs for Controlled Transformation Selection

### v1

@@ 6 Pairwise Comparison: WorldCollege Dataset
In this section, we apply our method to the World College dataset [^7] http://www.allourideas.org/worldcollege. This dataset contains 9,408 user-annotated pairwise comparisons among 261 colleges, collected via the All Our Ideas crowdsourcing platform; 8,544 of these comparisons were deemed valid. A diverse group of 400 annotators from various countries were asked to choose which university in a given pair they found more attractive to attend.
For our experiments, we subsample the dataset to exclude colleges with very few annotated comparisons. Our specific sub-dataset comprises $p=69$ colleges that rank in the top 80 of the QS rankings among the original 261 colleges. This subsampling yields $n=2112$ annotated pairwise comparisons among these 69 colleges, resulting in $m=608$ unique pairwise comparisons. More details on the QS rankings are in Supplementary Material Section I.
In this experiment, the Bradley-Terry model is adopted, and thus the transformation matrix $D\in\mathbb{R}^{m\times p}$ is taken as the graph difference operator of the subsampling graph, i.e. for each pair of colleges ($i_{1}<i_{2}$) compared by the annotator, there exists an unique row $j$ in $D$, such that $D_{j,i_{1}}=1$, $D_{j,i_{2}}=-1$, and $D_{j,k}=0$ for $k\neq i_{1},i_{2}$.
Figure 6: Selected Pairs of Colleges with Significant Preference Differences. Each point in the plot represent the pair of colleges with respective QS ranks marked by the column and row indices. To be specific, the dark blue points represent the pairs of colleges selected by our method to have significant preference differences, while the light blue points represent the pairs of colleges that are included in $D$. The diagonals are marked in deeper dark blue for better visualization.
Figure 6 presents the selection results for pairs of colleges with significant preference differences. As the figure shows, our method tends to select pairs located further from the diagonal with higher frequency than those closer to it. This indicates that our method effectively identifies university pairs with substantial QS ranking differences, which are more likely to represent true discoveries. For instance, our method identified the pairs Massachusetts Institute of Technology (MIT) (QS rank 1) and Osaka University (QS rank 63), and MIT (1) and Tohoku University (75), both exhibiting large ranking gaps. Other examples include Stanford University (2) and Lund University (73), and Harvard University (3) and Korea Advanced Institute of Science and Technology (47). These findings support the effectiveness of our approach in uncovering university pairs with notable differences in global academic standing.
@@
To further illustrate our method’s selection tendencies, Figure 7 presents a frequency plot displaying the selection frequencies of pairs against their QS ranking differences. The figure clearly shows an increasing trend in selection frequency as the QS ranking difference between pairs grows. This further validates our claim that our method tends to select university pairs with significant QS ranking differences and likely maintains a low False Discovery Rate (FDR). More specifically, our method consistently detects connections between colleges exhibiting substantial differences in their global rankings. For instance, connections between the Massachusetts Institute of Technology (rank 1) and Osaka University (rank 63), or Harvard University (rank 3) and Tohoku University (rank 75), were detected in 73 and 89 out of 100 experiments, respectively. The detected connections frequently involve pairs with rank differences exceeding 60, demonstrating our method’s particular sensitivity and effectiveness in uncovering relationships between institutions with large rank differences.
@@ 7 Conclusion
In this paper, we introduce the Model-X Split Knockoff method, an extension of both the Split Knockoff and Model-X Knockoff frameworks. This method effectively controls the false discovery rate in transformation selection problems across a broad range of non-linear statistical models. To overcome the inherent challenge posed by models combining a random design matrix with a deterministic linear transformation, we introduce a novel auxiliary randomized design. This randomization interacts with both the random design and the deterministic transformation, enabling the extraction of information related to transformational sparsity and successfully bypassing associated difficulties. Moreover, the flexibility of this auxiliary randomized design allows Model-X Split Knockoffs to achieve superior selection power compared to standard Model-X Knockoffs when both are applicable. Simulation experiments and real-world applications to Alzheimer’s Disease and the World College ranking validate the proposed method’s effectiveness.

### v2

@@ 6 Pairwise Comparison: WorldCollege Dataset
In this section, we apply our method to the World College dataset. For our experiments, we subsample the dataset to exclude colleges with very few annotated comparisons. Our specific sub-dataset comprises $p=69$ colleges that rank in the top 80 of the QS rankings among the original 261 colleges. This subsampling yields $n=2,112$ annotated pairwise comparisons among these 69 colleges, resulting in $m=608$ unique pairwise comparisons. More details on the descriptions of dataset and the QS rankings are in Supplementary Material Section I.
In this experiment, the Bradley-Terry model is adopted, and thus $D\in\mathbb{R}^{m\times p}$ is defined as: for each pair of college ($i_{1}<i_{2}$) compared by the annotator, there exists an unique row $j$ in $D$, such that $D_{j,i_{1}}=1$, $D_{j,i_{2}}=-1$, and $D_{j,k}=0$ for $k\neq i_{1},i_{2}$.
@@
Figure 3 presents the selection frequency from 100 runs of our method, showing a clear positive correlation with the QS ranking difference between pairs. This trend demonstrates that our method effectively identifies university pairs with significant ranking gaps, which suggests a low false discovery rate. For instance, connections between colleges with large rank differences, such as Massachusetts Institute of Technology (1) and Osaka University (63), or Harvard University (rank 3) and Tohoku University (rank 75), were detected in 73 and 89 out of 100 experiments respectively. The consistent selection of pairs with rank differences over 60 underscores the method’s sensitivity in uncovering substantial disparities.
@@ 7 Conclusion
We introduced the Model-X Split Knockoff method, a framework extending both Split and Model-X Knockoffs to control the FDR for transformation selection in a broad range of statistical models. Our key innovation, an auxiliary randomized design, resolves the challenge of combining random designs with deterministic transformations. This approach not only enables inference on transformational sparsity but also provides superior selection power over standard Model-X Knockoffs. The method’s effectiveness is validated by simulations and real-world applications.